OpenStudy (anonymous):

check for extraneous solutions. sqrt x+7-x=1 Please show step by step on how to do this. Thankyou.

OpenStudy (anonymous):

Let \(u=\sqrt{x}\), then x = \(u^2\) Now, the equation becomes \[u+7-u^2=1\] Can you do it?

OpenStudy (anonymous):

Wat?

OpenStudy (anonymous):

@Krazete

OpenStudy (anonymous):

I don't get the way you're explaining it to me @RolyPoly

OpenStudy (anonymous):

Please make use of parentheses. The way you write problems is sorta confusing. Is it √(x+7)-x=1 or ...?

OpenStudy (anonymous):

I guess. Sorry, it doesn't show parentheses on my math problem.

OpenStudy (anonymous):

Is the equation (1)\(\sqrt{ x+7-x} =1\) or (2) \(\sqrt{x}+7-x=1\)?

OpenStudy (anonymous):

nvm, i got the answer. Thx

OpenStudy (anonymous):

nvm.. I tried wolframalpha. But only get two step by steps each day:(

OpenStudy (anonymous):

I'll have to get the answer tomorrow.

OpenStudy (anonymous):

It would be better if you know how to solve it, instead of just relying on WolframAlpha.

OpenStudy (anonymous):

eh, true. But lets be honest with ourselves, I'm never going to use this advanced of math in the future.

OpenStudy (anonymous):

Honestly, I just need credits/passed classes

OpenStudy (anonymous):

you can argue all you want, but it's true.

OpenStudy (anonymous):

Don't want to argue, just want to help you pass the course and you can learn how to solve it at the same time :)

OpenStudy (anonymous):

can you just show me please? I'm 18 lessons behind:(

OpenStudy (nikato):

\[\sqrt{x+7}-x=1\]

OpenStudy (nikato):

add x to both sides

OpenStudy (anonymous):

It would be a different story if I wasn't that behind

OpenStudy (anonymous):

\[\sqrt{x-7-x}=1\]\[\sqrt{x-7}-x=1\]\[\sqrt{x}-7-x=1\] Which one is your question?

OpenStudy (anonymous):

like this @nikato